11. A pharmaceutical company makes pills that should contain 20.0 mg of the drug Phenobarbital. A random sample of 25 pills yielded a mean of 20.5 mg and standard deviation of 1.5 mg. At significance level001 test the claim that the pills are acceptable (H 20.0 mg). 603
Answer #1

Q.11) Given that, sample size ( n ) = 20

sample mean (ar x) = 20.5 mg

sample standard deviation ( s ) = 1.5 mg

significance level (alpha) = 0.03

The null and the alternative hypotheses are,

H0 : μ 20.0

Ha : μ 20,0

Test statistic is,

г. _μ 20.5-20 = 1.491 ˇ 丽 Vn

Degrees of freedom = n - 1 = 20 - 1= 19

p-value = 0.1524

Here, p-value = 0.1524 > Q0,03

So, we fail to reject the null hypothesis ( H0).